OpenStudy (anonymous):

points (3,1) and (1,4) lie on line k. what is the slope of the line that is perpendicular to k?

OpenStudy (anonymous):

What is the slope of line k in the first place? Rise / Run = slope or (y2- y1)/(x2-x1)= slope where x1 y1 x2 y2 come from (x1,y1) (x2,y2) These are the points given to you

OpenStudy (anonymous):

Slope of line K = (4-1)/(1-3) = -3/2 Slope of its perpendicular line would be +2/3 because the two slopes have to multiply up to -1.